Bryan Carpenter

Flat 8, Broughton House 28/30 Bullar Road Bitterne Park Southampton, S018 1GS

December 8, 1995

The Lettings Manager, London and Perth, 72 London Road, Southampton, SO15 2AJ

Dear sir,

I will shortly be leaving Southampton, and will therefore be vacating the property I rent through you—Flat 8, Broughton House.

I would like to terminate my tenancy one month from December 13th, i.e., on 13th January, 1996. I will probably move out of the property about a week before that date.

If you need to contact me, my daytime telephone number is (01703) 593368.

Yours sincerely,